### v1

# Leveraging LLMs for Enhanced Open-Vocabulary 3D Scene Understanding in Autonomous Driving

Abstract
This paper introduces a novel method for open-vocabulary 3D scene understanding in autonomous driving by combining Language Embedded 3D Gaussians with Large Language Models (LLMs) for enhanced inference. We propose utilizing LLMs to generate contextually relevant canonical phrases for segmentation and scene interpretation. Our method leverages the contextual and semantic capabilities of LLMs to produce a set of canonical phrases, which are then compared with the language features embedded in the 3D Gaussians. This LLM-guided approach significantly improves zero-shot scene understanding and detection of objects of interest, even in the most challenging or unfamiliar environments. Experimental results on the WayveScenes101 dataset demonstrate that our approach surpasses state-of-the-art methods in terms of accuracy and flexibility for open-vocabulary object detection and segmentation. This work represents a significant advancement towards more intelligent, context-aware autonomous driving systems, effectively bridging 3D scene representation with high-level semantic understanding.
keywords:
Semantic Scene Understanding, Language-Embedded3D Gaussians, LLM, Autonomous Driving
Figure 1: Overview of our LLM-enhanced Language Embedded 3D Gaussian Splatting pipeline. The LE3DGS model generates language feature maps, which are then processed using Algorithm 1 to compute relevancy scores with LLM-generated queries ($p_{\texttt{quer}}$), helping positives ($p_{\texttt{help}}$), and canonical phrases ($p_{\texttt{canon}}$). The system responds to high-level driving scenario queries by highlighting relevant objects (cars, traffic lights, pedestrians) across multiple views.
1 Introduction
Understanding complex scenes in three dimensions is crucial for advancements in autonomous systems, augmented reality, and robotics. The field of 3D scene understanding has seen significant advancements in recent years, particularly with the emergence of neural radiance fields (NeRF) [1] and 3D Gaussian Splatting (3DGS) [2] for novel view synthesis. However, the challenge of open-vocabulary querying in 3D space remains critical for tasks such as object localization and segmentation in autonomous driving scenarios.
Concurrently, large language models (LLMs) [3, 4, 5, 6, 7] have demonstrated remarkable abilities in grasping context and effectively analyzing and interpreting complex scenarios [8]. Recent works have begun to explore the integration of LLMs into autonomous driving tasks. For instance, NuPrompt [9], a large-scale dataset for 3D object detection and tracking in autonomous driving, utilizes natural language prompts to predict object trajectories across different views and frames. While this approach outperforms traditional methods, it primarily uses LLMs to enhance 3D detection tasks rather than directly influencing tracking [8]. Further advancements in this direction include HiLM-D [10], which leverages multimodal LLMs (MLLMs) [11, 12, 13] to improve autonomous driving tasks by integrating high-resolution information. Similarly, ToKEN [14] introduces a novel MM-LLM that tokenizes the world into object-level knowledge, enabling better utilization of LLM’s reasoning capabilities to enhance autonomous vehicle planning in long-tail scenarios.
Building upon these developments, our study proposes a novel method that combines LLMs with Language Embedded 3D Gaussians (LE3DGS) to improve autonomous vehicle scene understanding. This approach addresses the limitations of existing methods in open-vocabulary segmentation and retrieval. The key innovation lies in using LLMs to enhance text embeddings at inference with contextual understanding, enabling a more sophisticated and flexible perception system capable of interpreting complex environmental cues. Figure 1 illustrates the pipeline of our proposed approach.
By integrating the contextual reasoning capabilities of LLMs with the spatial precision of 3D Gaussians, our method aims to significantly improve the understanding of driving scenes. This integration not only enhances the system’s ability to identify and localize objects but also provides a richer semantic interpretation of the scene. The main contributions of our work are as follows:
  • We propose a novel integration of LLMs with LE3DGS for enhanced scene understanding in autonomous driving.
  • We demonstrate how LLMs can be used to generate contextually relevant queries and canonical phrases, improving the flexibility and adaptability of 3D scene analysis.
  • We present experimental results on the Wayvescene101 dataset, showcasing improved performance in both accuracy and adaptability of object detection and scene understanding compared to existing methods.
2 Related Work
Recent work has focused on incorporating semantic information into 3D scene representations. In [15], authors extended neural radiance fields (NeRF) to jointly encode semantics with appearance and geometry, enabling 2D semantic labels using a small amount of in-place annotations. Neural Feature Fusion Fields (N3F) [16] distills 2D image features into a 3D student network, improving dense 2D image feature extractors when applied to multiple images reconstructible as a 3D scene. Language Embedded Radiance Fields (LERF) [17] grounds language embeddings from models like CLIP [18] into NeRF, enabling open-ended language queries in 3D. This approach allows for pixel-aligned, zero-shot queries on distilled 3D CLIP embeddings without relying on region proposals or masks.
3D Gaussian Splatting [2] has emerged as an efficient method for high-quality novel view synthesis. It offers a rendering quality comparable to NeRF but with faster training and inference times [19]. Feature 3DGS [20] enables 3D Gaussian splatting on arbitrary-dimension semantic features via 2D foundation model distillation. This method showcases novel view semantic segmentation, language-guided editing, and segmenting anything through learning feature fields from state-of-the-art 2D foundation models. LE3DGS [21] addresses the challenge of embedding language features into 3D Gaussians without prohibitive memory usage. This approach uses a dedicated quantization scheme and a novel embedding procedure to achieve smooth yet accurate queries while countering multi-view feature inconsistencies. Authors in [22] presented ConceptFusion, an open-set multimodal 3D mapping approach that leverages foundation models pre-trained on internet-scale data to reason about concepts across modalities such as natural language, images, and audio. This method enables effective zero-shot spatial reasoning without additional training or fine-tuning. FeatureNeRF [23] learns generalizable NeRFs by distilling pre-trained vision foundation models. This approach allows mapping 2D images to continuous 3D semantic feature volumes, which can be used for various downstream tasks such as semantic key point transfer and object part segmentation.
Despite these advancements, methods such as LE3DGS [21], LERF [17], and LangSplat [24] rely on a fixed set of predefined canonical phrases during inference. This constraint limits their adaptability to novel or context-specific scenarios, particularly in the dynamic and unpredictable environment of autonomous driving. Moreover, CLIP’s text input length is significantly constrained due to its text encoder being limited to 77 tokens, primarily trained on short texts. This results in a reduced effective token length which has been shown that the actual usable length is just 20 tokens [25]. This limits the previous method’s inference capabilities to very short queries.
Our work addresses this limitation by introducing a novel approach that leverages LLMs to generate queries and canonical phrases at inference time. By doing so, we enable a more flexible and context-aware system that can adapt its language understanding to the specific nuances of each scene. This approach allows for a more accurate and comprehensive interpretation of the 3D environment, capturing subtle details and rare scenarios that might be overlooked by systems relying on predefined canonical language sets. Through this integration of LLMs with LE3DGS, we aim to significantly enhance the open-vocabulary capabilities and overall scene understanding in autonomous driving contexts.
3 Methodology
3.1 Scene Representation with 3D Gaussians
We use the LE3DGS approach [21] for generating 3D Gaussians with language features integrated. This method represents the scene using language embedded Gaussian points and allows for both open-vocabulary querying and high-quality novel view synthesis in an efficient manner. The method consists of several key components:
3.1.1 3D Gaussian Splatting
The method builds upon the efficient 3D Gaussian Splatting technique [2] for high-quality novel view synthesis, which represents the scene with 3D Gaussians initialized from a sparse point cloud generated via Structure-from-Motion (SfM). Each Gaussian is characterized by its position, anisotropic covariance, opacity, and color via spherical harmonics. Such a representation gives rise to an efficiently differentiable projection to 2D for real-time rendering. The final image derives from the alpha blending model used in volumetric rendering, represented as:
$C=\sum_{i=1}^{N}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))c_{i}$ (1)
where $C$ is the accumulated color along the viewing ray, $T_{i}$ is the accumulated transmittance to the $i$-th sample. $\sigma_{i}$ and $\delta_{i}$ denote the density and segment length of the $i$-th Gaussian, and $c_{i}$ is the color contribution of the $i$-th Gaussian.
3.1.2 Dense Language Feature Extraction
Dense language features are extracted from multi-view images using both CLIP [18] and DINO [26] models. These features can provide rich semantic information about the scene.
3.1.3 Feature Quantization
To reduce the high memory requirement of the raw language feature embeddings on dense 3D Gaussians, a quantization scheme [27] is proposed. It creates a discrete feature space, yielding semantic information compression without losing its effectiveness.
3.1.4 Compact Semantic Features on 3D Gaussians
Instead of directly embedding high-dimensional language features, LE3DGS learns compact semantic feature vectors for each 3D Gaussian. These are then decoded into discrete semantic indices using a small MLP.
3.1.5 Adaptive Spatial Smoothing
To handle semantic ambiguities arising from multi-view inconsistencies, LE3DGS introduces an adaptive smoothing mechanism [21]. The smoothing loss is:
$L_{smo}=\|s_{MLP}-s^{*}_{G}\|^{2}+\max(u^{*}_{G},w_{s})\|s^{*}_{MLP}-s_{G}\|^{2}$ (2)
where $s_{MLP}$ are smoothed semantic features generated by an MLP, $s_{G}$ represents the original semantic features of the 3D Gaussians, $u_{G}$ is the learned uncertainty value for each 3D Gaussian, and $w_{s}$ is a minimal weight for the semantic smoothing term. The * operator denotes stop gradient, preventing backpropagation through these terms.
3.1.6 Optimization
It performs a joint optimization over 3D Gaussians for appearance and semantic information. The total loss of the reconstruction and the semantics is:
$\mathcal{L}=\lambda_{s}\mathcal{L}_{s}+\lambda_{smo}\mathcal{L}_{smo}.$ (3)
$\mathcal{L}_{s}$ is the semantic loss:
$\mathcal{L}_{s}=\lambda_{CE}\mathcal{L}_{CE}+\lambda_{u}\mathcal{L}_{u}$ (4)
where $\mathcal{L}_{CE}$ is cross entropy loss and $\mathcal{L}_{u}$ is the uncertainty regularization loss. For further details, the reader is referred to [21].
3.1.7 Inference
During inference, the language feature map $F$ is computed as:
$\mathcal{F}=\mathcal{M}_{\text{infer }}\mathbf{S}$ (5)
where $\mathcal{M}_{\text{infer }}$ is the inferred distribution of semantic indices and $\mathbf{S}$ is the matrix of quantized language features. To detect objects in a 3D scene based on a given text prompt, the process for calculating a relevance map of $\mathcal{F}$ begins with computing the CLIP embedding $\phi_{\texttt{quer}}$ for the text query. Additionally, the CLIP embeddings $\phi^{i}_{\texttt{canon}}$ for a set of predefined canonical phrases is calculated. The final relevancy score for each $\phi_{lang}$ as presented in [17] is given by:
$\text{relevancy}=\min\left(\frac{\exp(\mathcal{F}\cdot\phi_{\texttt{quer}})}{\exp(\mathcal{F}\cdot\phi^{i}_{\texttt{canon}})+\exp(\mathcal{F}\cdot\phi_{\texttt{quer}})}\right)$ (6)
where $\cdot$ operation is cosine similarity between the two. This scoring mechanism measures how much the rendered embedding is closer to the query embedding with respect to the canonical embedding, indicating the most relevant elements of the scene for the given text prompt. All renderings in [21, 17, 24] use the same predefined canonical phrases: “object”, “things”, “stuff”, and “texture”.
3.2 Inference Using LLM
The canonical phrases serve as a baseline or reference point against which the query and rendered embeddings are compared. They help to contextualize the similarity between the query and the rendered embeddings. Ideally, the canonical phrases should be different from the query and represent a diverse set of concepts or objects that might appear in the scene. They should not be too similar to the query, as that would defeat their purpose.
We can determine if the rendered embedding is more similar to the query than to these baseline concepts by comparing the rendered embedding to both the query and the canonical phrases. If a rendered embedding is significantly nearer to the query than to any canonical phrase, it is probably highly relevant to the query. Conversely, if a rendered embedding is nearer to a canonical phrase than to the query, it likely has little relevance to the query. Accordingly, having a diverse set of canonical phrases helps to cover a wide range of potential scene elements, making the relevancy scoring more robust.
Therefore, we propose leveraging a LLM to generate contextually relevant queries and canonical phrases based on the initial scene description. This approach provides a more adaptive and comprehensive inference framework. We prompt the LLM to identify target objects in the scene, and it responds with:
  • A main query object,
  • Helping positive words: related terms or attributes of the main query, and
  • Canonical phrases: objects to differentiate from the query, including similar objects, nearby objects, and background elements.
This LLM-driven approach allows for dynamic adaptation to diverse driving scenarios, enhancing the system’s ability to understand and interpret complex scenes.
Algorithm 1 quantifies the relevance of language features to a given query, incorporating auxiliary information from helping queries and canonical phrases. It requires $\mathcal{M}$ the language feature map derived from the 3D Gaussians. LLM provides a primary query $p_{\texttt{quer}}$, a set of helping queries $p_{\texttt{help}}$ to provide context, and canonical phrases $p_{\texttt{canon}}$. The process begins by encoding the textual inputs using a CLIP text encoder in Lines 1 to 3. We concatenate query and helping query embeddings to the positive embeddings in Line 4. Then cosine similarities between positive embeddings and the language feature map and between canonical embeddings and the language feature map are computed in Lines 5 and 6. Finally, the relevancy score is calculated using a softmax-like formula in line 7. Notably, the relevancy score represents the probability that each pixel corresponds to the queried object. Subsequently, applying a threshold to these scores allows for the segmentation of the queried elements within the image.
Algorithm 1 Relevancy Score Calculation
0:  language feature map: $\mathcal{M}$, query, helping query, and canonical phrase from LLM: $p_{\texttt{quer}}$, $p_{\texttt{help}}$, $p_{\texttt{canon}}$
0:  relevancy score
1:  $\phi_{\texttt{quer}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{quer}})$
2:  $\phi_{\texttt{help}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{help}})$
3:  $\phi_{\texttt{canon}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{canon}})$
4:  $\phi_{\texttt{pos}}\leftarrow\text{CONCATENATE}(\phi_{\texttt{quer}},\phi_{\texttt{help}})$
5:  $\sigma_{\texttt{pos}}\leftarrow\text{COSINE SIMILARITY}(\phi_{\texttt{pos}},\mathcal{M})$
6:  $\sigma_{\texttt{canon}}\leftarrow\text{COSINE SIMILARITY}(\phi_{\texttt{canon}},\mathcal{M})$
7:  $\text{score}\leftarrow\min\left(\frac{\exp(\sigma_{\texttt{pos}})}{\exp(\sigma_{\texttt{canon}})+\exp(\sigma_{\texttt{pos}})}\right)$
8:  return  score
4 Experiments
4.1 Dataset
We use a subset of the WayveScenes101 dataset [28], designed specifically to advance novel view synthesis in challenging driving scenarios. We selected a subset of five representative scenes (IDs: 12, 21, 36, 78 and 88) totaling about 500 images. This selection balances computational feasibility with the need for diverse and challenging data, encompassing various environmental conditions, lighting variations, and complex urban environments. This allows us to thoroughly evaluate our approach in handling the complexities of autonomous driving environments. For quantitative evaluation of our method’s semantic understanding capabilities, we generated segmentation maps for the images in the evaluation set using a pre-trained Segformer model [29] on the Cityscapes dataset [30]. This semantic information allows us to assess semantic accuracy.
4.2 Implementation
We trained the LE3DGS using the same loss function as described in [21] on the Wayvesene101 dataset. We set $\lambda_{s}=0.5$, $\lambda_{CE}=\lambda_{smo}=\lambda_{u}=0.1$ and other weights to the deafault. The training process involved two main phases: dense semantic feature extraction and 3D Gaussian point optimization. For the semantic feature extraction, we set the number of epochs to 500, ensuring a rich semantic representation of the scene. In the optimization phase for 3D Gaussian points, we employed a 2D position gradient threshold of 0.0001, with a total of 20,000 iterations. The densification process was halted halfway through, at 10,000 iterations. All other hyperparameters were kept consistent with those used in [2]. The entire training process was conducted on a single RTX 4090 GPU, taking approximately 40 minutes to complete for each scene in the dataset. The mean peak signal-to-noise ratio (PSNR) achieved is around 25.14.
Table 1: Input structure for GPT-3.5 Turbo query generation in our implementation. This setup guides the language model to provide structured information about objects relevant to driving scenarios.
Role Content
System A helpful assistant for a computer vision task, providing structured information about objects to pay attention to while driving. Include both important objects and nearby objects that might be at the borders. Instructions for formatting the response, including: Main Positive (object to pay attention to) 1-4 Helping Positives (related terms or attributes) 4-6 Negatives (objects to differentiate from, including similar objects, nearby objects, and background elements)
User "Driving through an intersection in an {Road Type} on a {Weather} {Time of Day}, what objects should the driver pay attention to? " or show the {Object}. "
For the inference phase, we leveraged the API of GPT-3.5 Turbo [31] to dynamically generate queries. For each primary query, the model also provided 1 to 4 helping positive words and 4 to 6 canonical phrases. Table 1 denotes the input structure to the LLM model. The {Road Type}, {Weather}, and {Time of Day} in the user prompts could be provided from the dataset scene metadata [28]. We then calculated the relevancy score using Algorithm 1. To obtain object segmentations, we applied a threshold of 0.5 to the computed relevancy scores. Figure 1 illustrates our inference approach pipeline for scene 88 in Wayvescene101 dataset. The canonical phrases used in the baseline approach are “object”, “things”, “stuff”, and “texture”, consistent with the approach described in [21].
Reconstructed
Predefined $p_{\texttt{canon}}$
Ours w/o $p_{\texttt{help}}$
Ours
Figure 2: Visual comparison of query relevance across different scenes with different approaches. From top to bottom query words are: “trees”, “traffic signs”, “pedestrian”, “sidewalk”, and “cars”.
4.3 Results
Our approach demonstrates significant improvements in open-vocabulary segmentation across various urban driving scenarios. Figure 2 provides a visual comparison, showcasing the progression from overly broad segmentation using predefined canonical phrases to refined results with our full LLM-enhanced method. Our approach accurately identifies and segments various elements such as trees, traffic lights, and sidewalks with high precision, illustrating the benefits of LLM-generated context-aware queries and helping positives.
Inference Method $\mathrm{Accuracy}\uparrow$ $\mathrm{Presicion}\uparrow$ $\mathrm{mIoU}\uparrow$ $\mathrm{mAP}\uparrow$
Predefined Canonical Phrase 0.65 0.33 0.23 0.49
Ours w/o $p_{\texttt{help}}$ 0.84 0.42 0.33 0.53
Ours 0.92 0.58 0.37 0.54
Table 2: Comparison of inference methods: our approach (with and without helping positive words $p_{\texttt{help}}$) versus predefined canonical phrases.
Quantitative results in Table 2 support these visual observations. Our full method achieves the highest performance across all metrics, with notable improvements in Accuracy (0.92), Precision (0.58), and mean Intersection over Union (mIoU) (0.37). These represent significant gains over the predefined canonical phrase method (0.65, 0.33, and 0.23 respectively). The ablation study (Ours w/o $p_{\texttt{help}}$) indicates the robustness of our LLM-driven query generation, outperforming the baseline even without helping positive words.
(a) Mean Accuracy
(b) Mean IoU
(c) Mean Precision
Figure 3: Performance comparison of our methods against baseline across different object classes for the scenes in dataset.
Figure 3 further breaks down performance across various urban scene elements. Our method shows significant improvements in Mean IoU and Mean Precision for challenging classes like "sidewalk", "traffic light", and "traffic signs", while maintaining an edge in common classes such as "building" and "car". These results collectively demonstrate the effectiveness of our LLM-enhanced approach in improving scene understanding and object segmentation across diverse urban environments.
For more detailed qualitative results, please refer to section A.1 in the appendix.
5 Discussion
Our approach demonstrates significant improvements in open-vocabulary 3D scene understanding, but there are several important aspects to consider:
Limitations of 3D Gaussian Splatting
While 3D Gaussian Splatting is effective for certain scenes, it struggles with generating large and dynamic scenes [32]. This limitation can affect the overall performance and applicability of our approach in more dynamic environments, such as busy urban intersections or highways with fast-moving vehicles.
System Integration and Real-time Performance
The integration of our approach with existing autonomous driving systems presents both opportunities and challenges. While our method offers enhanced semantic understanding, its real-time performance in the context of rapidly changing driving scenarios requires further optimization. Current autonomous systems often rely on faster, albeit less semantically rich, perception methods. Bridging this gap involves not only improving our method’s computational efficiency but also developing strategies for seamless integration with existing sensor fusion and decision-making pipelines.
6 Conclusion
In this study, we proposed a novel method that integrates LLMs with LE3DGS to improve autonomous vehicle scene understanding. By leveraging the robust reasoning capabilities of LLMs and the detailed spatial representation of 3D Gaussians, our approach significantly improves the accuracy and precision of open vocabulary object segmentation and scene interpretation tasks. Our experimental results indicate that our method outperforms traditional predefined canonical phrase-based approaches on all aspects of accuracy, precision, and mean Intersection over Union.
Future work will focus on enhancing our system’s capability to handle dynamic scenes by integrating LLMs with advanced 3D Gaussian Splatting techniques. We aim to fuse our approach with multi-modal sensor data, such as LiDAR, to create a more robust and comprehensive perception system. Another crucial next step is to seamlessly integrate our advanced semantic understanding with autonomous driving planners. This integration will involve developing efficient interfaces that translate our rich semantic insights into actionable information, directly informing path planning and decision-making algorithms for safer and more intelligent autonomous navigation.
Appendix A Appendix
A.1 Qualitative Results
Figures 4, 5, and 6 showcase our approach’s effectiveness in open vocabulary segmentation across diverse driving scenes. Each figure presents, from top to bottom: ground truth images, Segformer-generated segmentation maps (used for quantitative evaluation), our high-quality scene reconstructions using LE3DGS, and our open vocabulary segmentation results. Our method accurately highlights key elements like cars, pedestrians, trees, buildings, and traffic signs/lights. For quantitative evaluation, we align with Segformer classes [33], but our visualizations include additional classes to demonstrate the flexibility of our approach. These figures illustrate our method’s capability to accurately reconstruct and semantically segment complex urban scenes, even for classes beyond traditional segmentation models.
Figure 4: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scene 88.
(a) Scene 12
(b) Scene 21
Figure 5: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scenes 12 and 21.
(a) Scene 36
(b) Scene 78
Figure 6: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scenes 36 and 78.

### v3

@@
-Leveraging LLMs for Enhanced Open-Vocabulary 3D Scene Understanding in Autonomous Driving
+Query3D: LLM-Powered Open-Vocabulary Scene Segmentation with Language Embedded 3D Gaussians
 Abstract
-This paper introduces a novel method for open-vocabulary 3D scene understanding in autonomous driving by combining Language Embedded 3D Gaussians with Large Language Models (LLMs) for enhanced inference.
-We propose utilizing LLMs to generate contextually relevant canonical phrases for segmentation and scene interpretation.
-Our method leverages the contextual and semantic capabilities of LLMs to produce a set of canonical phrases, which are then compared with the language features embedded in the 3D Gaussians.
-This LLM-guided approach significantly improves zero-shot scene understanding and detection of objects of interest, even in the most challenging or unfamiliar environments.
-Experimental results on the WayveScenes101 dataset demonstrate that our approach surpasses state-of-the-art methods in terms of accuracy and flexibility for open-vocabulary object detection and segmentation.
-This work represents a significant advancement towards more intelligent, context-aware autonomous driving systems, effectively bridging 3D scene representation with high-level semantic understanding.
-keywords:
-Semantic Scene Understanding, Language-Embedded3D Gaussians, LLM, Autonomous Driving
-Figure 1: Overview of our LLM-enhanced Language Embedded 3D Gaussian Splatting pipeline.
-The LE3DGS model generates language feature maps, which are then processed using Algorithm 1 to compute relevancy scores with LLM-generated queries ($p_{\texttt{quer}}$), helping positives ($p_{\texttt{help}}$), and canonical phrases ($p_{\texttt{canon}}$).
-The system responds to high-level driving scenario queries by highlighting relevant objects (cars, traffic lights, pedestrians) across multiple views.
+This paper introduces a novel method for open-vocabulary 3D scene querying in autonomous driving by combining Language Embedded 3D Gaussians with Large Language Models (LLMs).
+We propose utilizing LLMs to generate both contextually canonical phrases and helping positive words for enhanced segmentation and scene interpretation.
+Our method leverages GPT-3.5 Turbo as an expert model to create a high-quality text dataset, which we then use to fine-tune smaller, more efficient LLMs for on-device deployment.
+Our comprehensive evaluation on the WayveScenes101 dataset demonstrates that LLM-guided segmentation significantly outperforms traditional approaches based on predefined canonical phrases.
+Notably, our fine-tuned smaller models achieve performance comparable to larger expert models while maintaining faster inference times.
+Through ablation studies, we discover that the effectiveness of helping positive words correlates with model scale, with larger models better equipped to leverage additional semantic information.
+This work represents a significant advancement towards more efficient, context-aware autonomous driving systems, effectively bridging 3D scene representation with high-level semantic querying while maintaining practical deployment considerations.
+Code and additional resources are available at https://github.com/Zhourobotics/Query-3DGS-LLM.
 1 Introduction
@@
 However, the challenge of open-vocabulary querying in 3D space remains critical for tasks such as object localization and segmentation in autonomous driving scenarios.
+Figure 1: Overview of our LLM-enhanced Language Embedded 3D Gaussian Splatting pipeline.
+The LE3DGS model generates language feature maps, which are then processed using Algorithm 1 to compute relevancy scores with LLM-generated queries ($p_{\texttt{quer}}$), helping positives ($p_{\texttt{help}}$), and canonical phrases ($p_{\texttt{canon}}$).
+The system responds to the high-level queries of a driving scenario by highlighting relevant objects (cars, traffic lights, pedestrians) across multiple views.
 Concurrently, large language models (LLMs) [2, 1, 31, 15, 11] have demonstrated remarkable abilities in grasping context and effectively analyzing and interpreting complex scenarios [39].
@@
 Further advancements in this direction include HiLM-D [8], which leverages multimodal LLMs (MLLMs) [45, 18, 23] to improve autonomous driving tasks by integrating high-resolution information.
-Similarly, ToKEN [14] introduces a novel MM-LLM that tokenizes the world into object-level knowledge, enabling better utilization of LLM’s reasoning capabilities to enhance autonomous vehicle planning in long-tail scenarios.
-Building upon these developments, our study proposes a novel method that combines LLMs with Language Embedded 3D Gaussians (LE3DGS) to improve autonomous vehicle scene understanding.
-This approach addresses the limitations of existing methods in open-vocabulary segmentation and retrieval.
-The key innovation lies in using LLMs to enhance text embeddings at inference with contextual understanding, enabling a more sophisticated and flexible perception system capable of interpreting complex environmental cues.
-Figure 1 illustrates the pipeline of our proposed approach.
-By integrating the contextual reasoning capabilities of LLMs with the spatial precision of 3D Gaussians, our method aims to significantly improve the understanding of driving scenes.
-This integration not only enhances the system’s ability to identify and localize objects but also provides a richer semantic interpretation of the scene.
+Similarly, ToKEN [30] introduces a novel MLLM that tokenizes the world into object-level knowledge, enabling better utilization of LLM’s reasoning capabilities.
+Recent works like Language Embedded Radiance Fields (LERF) [20] and Language Embedded 3D Gaussians (LE3DGS) [28] have demonstrated the potential of language-embedded neural and 3D gaussians representations for scene understanding.
+Our study extends these approaches by introducing LLM-powered query processing to enhance open-vocabulary querying and segmentation for the scene understanding of autonomous vehicles.
+Our system supports both explicit queries that directly specify objects of interest (e.g., “Show me the pedestrians”) and contextual queries that require scene understanding (e.g., “What should I watch for at this intersection?”).
+For either type of query, our approach employs LLMs to dynamically generate three complementary types of semantic references: query words that capture the core search intent, context-aware canonical phrases that provide semantic anchors, and helping positive words that enhance the segmentation further.
+Figure 1 demonstrates our approach processing a contextual query.
+The LLM first interprets the user’s query to identify relevant scene elements, then generates appropriate canonical phrases and helping positive words.
+This combination enables more precise object segmentation than traditional fixed-phrase methods, particularly for contextual queries that require deeper scene understanding.
+By integrating the contextual reasoning capabilities of LLMs with the spatial precision of 3D Gaussians, our method significantly improves the querying capabilities in driving scenes.
+This integration enables more flexible and context-aware scene interpretation while providing an efficient path toward practical deployment.
 The main contributions of our work are as follows:
-  • We propose a novel integration of LLMs with LE3DGS for enhanced scene understanding in autonomous driving.
-  • We demonstrate how LLMs can be used to generate contextually relevant queries and canonical phrases, improving the flexibility and adaptability of 3D scene analysis.
-  • We present experimental results on the Wayvescene101 dataset, showcasing improved performance in both accuracy and adaptability of object detection and scene understanding compared to existing methods.
+  • We enhance LE3DGS with LLM-generated language guidance.
+Unlike previous approaches that use fixed canonical phrases, our LLM dynamically generates two types of descriptive language: canonical phrases to differentiate scene elements and helping positive words for additional context.
+These LLM-generated phrases are more contextually suited for LE3DGS’s language embeddings, leading to more accurate scene segmentation.
+This is particularly effective for complex queries where traditional fixed phrases might miss subtle distinctions.
+  • We develop an effective fine-tuning strategy that enables smaller LLMs to achieve performance comparable to larger expert models while maintaining faster inference times for on-device deployment in autonomous vehicles.
+  • Through comprehensive experiments on the WayveScene101 dataset, we demonstrate that our LLM-guided approach significantly outperforms traditional fixed canonical phrase methods.
+Our ablation studies reveal that the effectiveness of helping positive words correlates with model scale, with larger models better equipped to leverage additional semantic context.
 2 Related Work
-Recent work has focused on incorporating semantic information into 3D scene representations.
-In [15], authors extended neural radiance fields (NeRF) to jointly encode semantics with appearance and geometry, enabling 2D semantic labels using a small amount of in-place annotations.
-Neural Feature Fusion Fields (N3F) [16] distills 2D image features into a 3D student network, improving dense 2D image feature extractors when applied to multiple images reconstructible as a 3D scene.
-Language Embedded Radiance Fields (LERF) [17] grounds language embeddings from models like CLIP [18] into NeRF, enabling open-ended language queries in 3D.
-This approach allows for pixel-aligned, zero-shot queries on distilled 3D CLIP embeddings without relying on region proposals or masks.
-3D Gaussian Splatting [2] has emerged as an efficient method for high-quality novel view synthesis.
-It offers a rendering quality comparable to NeRF but with faster training and inference times [19].
-Feature 3DGS [20] enables 3D Gaussian splatting on arbitrary-dimension semantic features via 2D foundation model distillation.
-This method showcases novel view semantic segmentation, language-guided editing, and segmenting anything through learning feature fields from state-of-the-art 2D foundation models.
-LE3DGS [21] addresses the challenge of embedding language features into 3D Gaussians without prohibitive memory usage.
-This approach uses a dedicated quantization scheme and a novel embedding procedure to achieve smooth yet accurate queries while countering multi-view feature inconsistencies.
-Authors in [22] presented ConceptFusion, an open-set multimodal 3D mapping approach that leverages foundation models pre-trained on internet-scale data to reason about concepts across modalities such as natural language, images, and audio.
-This method enables effective zero-shot spatial reasoning without additional training or fine-tuning.
-FeatureNeRF [23] learns generalizable NeRFs by distilling pre-trained vision foundation models.
-This approach allows mapping 2D images to continuous 3D semantic feature volumes, which can be used for various downstream tasks such as semantic key point transfer and object part segmentation.
-Despite these advancements, methods such as LE3DGS [21], LERF [17], and LangSplat [24] rely on a fixed set of predefined canonical phrases during inference.
-This constraint limits their adaptability to novel or context-specific scenarios, particularly in the dynamic and unpredictable environment of autonomous driving.
-Moreover, CLIP’s text input length is significantly constrained due to its text encoder being limited to 77 tokens, primarily trained on short texts.
-This results in a reduced effective token length which has been shown that the actual usable length is just 20 tokens [25].
-This limits the previous method’s inference capabilities to very short queries.
-Our work addresses this limitation by introducing a novel approach that leverages LLMs to generate queries and canonical phrases at inference time.
-By doing so, we enable a more flexible and context-aware system that can adapt its language understanding to the specific nuances of each scene.
-This approach allows for a more accurate and comprehensive interpretation of the 3D environment, capturing subtle details and rare scenarios that might be overlooked by systems relying on predefined canonical language sets.
-Through this integration of LLMs with LE3DGS, we aim to significantly enhance the open-vocabulary capabilities and overall scene understanding in autonomous driving contexts.
+2.1 3D Scene Representations with Semantic Features
+Recent advances in 3D scene representation have focused on incorporating semantic information alongside geometry and appearance.
+Neural radiance fields were extended to encode semantics [43], enabling 2D semantic labels with minimal annotations, while Neural Feature Fusion Fields (N3F) [32] improved feature extraction through 2D-to-3D distillation.
+3D Gaussian Splatting [19] emerged as an efficient alternative to NeRF, offering comparable quality with faster processing [22].
+The integration of language features into these representations marked a significant advancement.
+Language Embedded Radiance Fields (LERF) [20] pioneered open-ended language queries by grounding CLIP [27] embeddings into NeRF.
+Feature 3DGS [44] and Language Embedded 3D Gaussians (LE3DGS) [28] extended this concept to Gaussian splatting, enabling efficient language-guided editing and semantic segmentation.
+ConceptFusion [16] and FeatureNeRF [40] further advanced cross-modal reasoning by leveraging foundation models.
+2.2 LLM-based 3D Object Grounding
+Recent works have explored LLMs for 3D object grounding tasks with different approaches [12, 3, 17, 9].
+ViL3DRel [5] incorporates spatial relationships through a transformer-based architecture with spatial self-attention.
+LLM-Grounder [38] and VLM-Grounder [36] employ LLMs as reasoning agents for grounding, with the former decomposing complex queries into subtasks and the latter focusing on efficient multi-view processing.
+Object-centric approaches [14, 21] represent scenes through distinct object entities with unique identifiers, combining 2D and 3D features for comprehensive scene understanding.
+While these methods excel at grounding specific objects based on language descriptions, our work differs by focusing on enhancing open-vocabulary querying capabilities through LLM-generated canonical phrases and helping positives, enabling more flexible scene interpretation for autonomous driving applications.
+2.3 Limitations of Current Approaches
+Despite these advances, current methods including LE3DGS [28], LERF [20], and LangSplat [26] share several key limitations.
+They rely on fixed predefined canonical phrases during inference, which limits their adaptability to both explicit and contextual queries.
+Their underlying vision-language models exhibit “bag-of-words” behavior [41], treating queries as collections of independent words rather than understanding their semantic relationships.
+Additionally, the reliance on CLIP’s text encoder, with its effective usable length of just 20 tokens [42], restricts these methods to very short queries.
+Our work addresses these limitations by leveraging LLMs to dynamically generate canonical phrases and helping positive words at inference time.
+By integrating LLMs with LE3DGS, we enable a system that can effectively process both explicit and contextual queries while adapting its language understanding to specific scene requirements, capturing subtle details that the systems with fixed canonical phrases might miss.
+Figure 2: Visual comparison of segmentation results using different Qwen models responding to a contextual query “Driving through an intersection in an urban area on a sunny day, what objects should the driver pay attention to?” From top to bottom: Results from Qwen-0.5B, 1.5B, 3B, and 7B models.
+Each row shows three camera views (left, front, right) of the same scene.
+The results demonstrate how larger models produce more precise and contextually relevant segmentations for autonomous driving scenarios.
 3 Methodology
 3.1 Scene Representation with 3D Gaussians
-We use the LE3DGS approach [21] for generating 3D Gaussians with language features integrated.
-This method represents the scene using language embedded Gaussian points and allows for both open-vocabulary querying and high-quality novel view synthesis in an efficient manner.
-The method consists of several key components:
+While our method is compatible with various neural representations like LERF [20], we leverage LE3DGS [28] for its efficient inference capabilities to generate 3D Gaussians with language features integrated.
+This method represents the scene using language-embedded Gaussian points and allows for both open-vocabulary querying and high-quality novel view synthesis in an efficient manner.
+The method consists of several key components as follows:
 3.1.1 3D Gaussian Splatting
@@
 where $\mathcal{L}_{CE}$ is cross entropy loss and $\mathcal{L}_{u}$ is the uncertainty regularization loss. For further details, the reader is referred to [28].
-3.1.7 Inference
-During inference, the language feature map $F$ is computed as:
+3.2 Inference
+During inference, we compute a language feature map $\mathcal{F}$ from the scene representation:
 $\mathcal{F}=\mathcal{M}_{\text{infer }}\mathbf{S}$ (5)
 where $\mathcal{M}_{\text{infer }}$ is the inferred distribution of semantic indices and $\mathbf{S}$ is the matrix of quantized language features.
-To detect objects in a 3D scene based on a given text prompt, the process for calculating a relevance map of $\mathcal{F}$ begins with computing the CLIP embedding $\phi_{\texttt{quer}}$ for the text query.
-Additionally, the CLIP embeddings $\phi^{i}_{\texttt{canon}}$ for a set of predefined canonical phrases is calculated.
-The final relevancy score for each $\phi_{lang}$ as presented in [17] is given by:
-$\text{relevancy}=\min\left(\frac{\exp(\mathcal{F}\cdot\phi_{\texttt{quer}})}{\exp(\mathcal{F}\cdot\phi^{i}_{\texttt{canon}})+\exp(\mathcal{F}\cdot\phi_{\texttt{quer}})}\right)$ (6)
-where $\cdot$ operation is cosine similarity between the two.
-This scoring mechanism measures how much the rendered embedding is closer to the query embedding with respect to the canonical embedding, indicating the most relevant elements of the scene for the given text prompt.
-All renderings in [21, 17, 24] use the same predefined canonical phrases: “object”, “things”, “stuff”, and “texture”.
-3.2 Inference Using LLM
-The canonical phrases serve as a baseline or reference point against which the query and rendered embeddings are compared.
-They help to contextualize the similarity between the query and the rendered embeddings.
-Ideally, the canonical phrases should be different from the query and represent a diverse set of concepts or objects that might appear in the scene.
-They should not be too similar to the query, as that would defeat their purpose.
-We can determine if the rendered embedding is more similar to the query than to these baseline concepts by comparing the rendered embedding to both the query and the canonical phrases.
-If a rendered embedding is significantly nearer to the query than to any canonical phrase, it is probably highly relevant to the query.
-Conversely, if a rendered embedding is nearer to a canonical phrase than to the query, it likely has little relevance to the query.
-Accordingly, having a diverse set of canonical phrases helps to cover a wide range of potential scene elements, making the relevancy scoring more robust.
-Therefore, we propose leveraging a LLM to generate contextually relevant queries and canonical phrases based on the initial scene description.
-This approach provides a more adaptive and comprehensive inference framework.
-We prompt the LLM to identify target objects in the scene, and it responds with:
-  • A main query object,
-  • Helping positive words: related terms or attributes of the main query, and
-  • Canonical phrases: objects to differentiate from the query, including similar objects, nearby objects, and background elements.
-This LLM-driven approach allows for dynamic adaptation to diverse driving scenarios, enhancing the system’s ability to understand and interpret complex scenes.
-Algorithm 1 quantifies the relevance of language features to a given query, incorporating auxiliary information from helping queries and canonical phrases.
-It requires $\mathcal{M}$ the language feature map derived from the 3D Gaussians.
-LLM provides a primary query $p_{\texttt{quer}}$, a set of helping queries $p_{\texttt{help}}$ to provide context, and canonical phrases $p_{\texttt{canon}}$.
-The process begins by encoding the textual inputs using a CLIP text encoder in Lines 1 to 3.
-We concatenate query and helping query embeddings to the positive embeddings in Line 4.
-Then cosine similarities between positive embeddings and the language feature map and between canonical embeddings and the language feature map are computed in Lines 5 and 6.
-Finally, the relevancy score is calculated using a softmax-like formula in line 7.
-Notably, the relevancy score represents the probability that each pixel corresponds to the queried object.
-Subsequently, applying a threshold to these scores allows for the segmentation of the queried elements within the image.
-Algorithm 1 Relevancy Score Calculation
-0:  language feature map: $\mathcal{M}$, query, helping query, and canonical phrase from LLM: $p_{\texttt{quer}}$, $p_{\texttt{help}}$, $p_{\texttt{canon}}$
+The relevancy score for scene elements is calculated following [20]:
+$\text{relevancy}=\min_{i}\frac{\exp(\mathcal{F}\cdot\phi_{\texttt{quer}})}{\exp(\mathcal{F}\cdot\phi^{i}_{\texttt{canon}})+\exp(\mathcal{F}\cdot\phi_{\texttt{quer}})}$ (6)
+where $\phi_{\texttt{quer}}$ is the CLIP embedding of the query text, $\phi^{i}_{\texttt{canon}}$ are embeddings of canonical phrases, and $\cdot$ represents cosine similarity.
+3.3 LLM-Enhanced Open-Vocabulary Querying
+Previous approaches [28, 20, 26] rely on fixed canonical phrases (“object”, “things”, “stuff”, “texture”), limiting their ability to handle diverse scenarios.
+Additionally, these methods often exhibit the “bag-of-words” behavior [38], treating queries as collections of independent words rather than understanding their semantic relationships.
+Our system processes both explicit queries that directly specify objects (e.g., “Show me the traffic signs”) and contextual queries that require scene interpretation (e.g., “What should I watch for in this construction zone?”).
+For either query type, our LLM generates:
+  • Canonical phrases for computing relevancy scores to differentiate scene elements.
+  • Helping positive words that enhance matching with LE3DGS’s language embeddings.
+For example, when processing either an explicit query “Show me the traffic signs” or interpreting a contextual query about a construction zone, our LLM generates:
+  • Canonical phrases to differentiate elements: “construction cones”, “traffic cones”, “pedestrians”, “vehicles”, “traffic lights”, and
+  • Helping positive words for better matching: “road signs”, “warning signs”, “directional markers”.
+The rationale for this approach is twofold.
+First, unlike CLIP’s limited token capacity [42], LLMs can process longer, more nuanced descriptions to generate appropriate phrases.
+Second, LLMs understand semantic relationships between concepts, addressing the bag-of-words limitation [41, 20] by providing related terms that help distinguish fine-grained differences.
+This LLM-driven approach enables more accurate segmentation by adapting to both explicit and contextual queries while maintaining precise object differentiation through canonical phrases and helping positive words.
+Its detailed steps are shown in Algorithm 1
+Algorithm 1 Query Processing and Relevancy Score Calculation
+0:  language feature map: $\mathcal{F}$, user query (explicit or contextual)
 0:  relevancy score
-1:  $\phi_{\texttt{quer}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{quer}})$
-2:  $\phi_{\texttt{help}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{help}})$
-3:  $\phi_{\texttt{canon}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{canon}})$
-4:  $\phi_{\texttt{pos}}\leftarrow\text{CONCATENATE}(\phi_{\texttt{quer}},\phi_{\texttt{help}})$
-5:  $\sigma_{\texttt{pos}}\leftarrow\text{COSINE SIMILARITY}(\phi_{\texttt{pos}},\mathcal{M})$
-6:  $\sigma_{\texttt{canon}}\leftarrow\text{COSINE SIMILARITY}(\phi_{\texttt{canon}},\mathcal{M})$
-7:  $\text{score}\leftarrow\min\left(\frac{\exp(\sigma_{\texttt{pos}})}{\exp(\sigma_{\texttt{canon}})+\exp(\sigma_{\texttt{pos}})}\right)$
-8:  return  score
+1:  $p_{\texttt{quer}},p_{\texttt{help-pos}},p_{\texttt{canon}}\leftarrow\text{LLM}(\text{query})$
+2:  $\phi_{\texttt{quer}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{quer}})$
+3:  $\phi_{\texttt{help-pos}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{help-pos}})$
+4:  $\phi_{\texttt{canon}}\leftarrow\text{CLIP}_{\text{text encoder}}(p_{\texttt{canon}})$
+5:  $\phi_{\texttt{pos}}\leftarrow\text{CONCATENATE}(\phi_{\texttt{quer}},\phi_{\texttt{help-pos}})$
+6:  $\sigma_{\texttt{pos}}\leftarrow\text{COSINE SIMILARITY}(\phi_{\texttt{pos}},\mathcal{F})$
+7:  $\sigma_{\texttt{canon}}\leftarrow\text{COSINE SIMILARITY}(\phi_{\texttt{canon}},\mathcal{F})$
+8:  $\text{score}\leftarrow\min_{i}\left(\frac{\exp(\sigma_{\texttt{pos}})}{\exp(\sigma_{\texttt{canon}}^{i})+\exp(\sigma_{\texttt{pos}})}\right)$
+9:  return  score
+Algorithm 1 processes user queries and calculates relevancy scores for scene elements.
+The algorithm accepts both explicit queries (e.g., “Show me pedestrians”) and contextual queries (e.g., “What should I watch at this intersection?”).
+In Line 1, the LLM processes the user query to generate three outputs: primary query $p_{\texttt{quer}}$ identifying objects of interest, helping positive words $p_{\texttt{help-pos}}$ for enhanced matching, and canonical phrases $p_{\texttt{canon}}$ for differentiation.
+In Lines 24, these phrases are encoded using CLIP’s text encoder to obtain their respective embeddings.
+Line 5 concatenates the primary query and helping positive embeddings into $\phi_{\texttt{pos}}$ to enhance matching accuracy.
+Lines 67 compute cosine similarities between the language feature map and both the combined positive embeddings ($\sigma_{\texttt{pos}}$) and canonical embeddings ($\sigma_{\texttt{canon}}$).
+Finally, Line 8 applies the softmax-like formula shown in Eq.
+6 to calculate the relevancy score, representing the probability of each pixel corresponding to the queried objects.
+The resulting scores can be thresholded to obtain the final segmentation.
+System: Object detection assistant Scene: “Driving through an urban street with construction work, traffic cones, and barriers on an overcast day.
+The road has temporary lane markings and directional signs, with buildings on both sides and leafless trees along the street.” Task: For a given object: Provide up to 3 helping positives, and Provide up to 6 Canonical phrases.
+Query: Object: “Pedestrian” Response: Helping positives: person, human, crosswalk
+Canonical phrases: traffic light, road, traffic cone, building, tree
+Figure 3: Example conversation with GPT-3.5 Turbo to query “Pedestrian” from scene 12 in WayveScene101.
+Table 1: Performance comparison of different model variants with and without helping positives.
+The table shows metrics for both instruction-tuned and fine-tuned Qwen and Llama models of varying sizes (0.5B-7B parameters) and GPT-3.5 Turbo against baseline segmentation.
+The best values in each column are in bold, second-best values are underlined.
+Model Without Helping Positives With Helping Positives
+IoU Accuracy Precision mAP IoU Accuracy Precision mAP
+Qwen2.5-0.5B-Instruct 0.16 0.68 0.22 0.42 0.08 0.75 0.13 0.32
+Qwen2.5-1.5B-Instruct 0.19 0.78 0.23 0.46 0.07 0.88 0.10 0.38
+Qwen2.5-3B-Instruct 0.18 0.75 0.24 0.47 0.05 0.77 0.15 0.34
+Qwen2.5-7B-Instruct 0.25 0.83 0.32 0.53 0.15 0.88 0.24 0.47
+Qwen2.5-0.5B-Finetuned 0.17 0.86 0.24 0.51 0.15 0.87 0.22 0.42
+Qwen2.5-1.5B-Finetuned 0.28 0.81 0.37 0.54 0.17 0.81 0.28 0.43
+Qwen2.5-3B-Finetuned 0.27 0.86 0.35 0.56 0.17 0.86 0.24 0.45
+Qwen2.5-7B-Finetuned 0.24 0.86 0.33 0.53 0.23 0.89 0.34 0.47
+Llama3.2-1B-Instruct 0.18 0.71 0.20 0.46 0.16 0.73 0.19 0.42
+Llama3.2-3B-Instruct 0.21 0.79 0.27 0.51 0.12 0.86 0.24 0.41
+Llama3.2-1B-Finetuned 0.18 0.76 0.24 0.49 0.12 0.81 0.22 0.40
+Llama3.2-3B-Finetuned 0.26 0.84 0.32 0.53 0.18 0.87 0.26 0.43
+GPT-3.5 Turbo 0.27 0.90 0.37 0.54 0.32 0.94 0.50 0.54
+Base 0.18 0.64 0.19 0.50 0.18 0.64 0.19 0.50
 4 Experiments
 4.1 Dataset
-We use a subset of the WayveScenes101 dataset [28], designed specifically to advance novel view synthesis in challenging driving scenarios.
+4.1.1 WayveScene
+We use a subset of the WayveScenes101 dataset [46], designed specifically to advance novel view synthesis in challenging real-world driving scenarios.
 We selected a subset of five representative scenes (IDs: 12, 21, 36, 78 and 88) totaling about 500 images.
@@
 This allows us to thoroughly evaluate our approach in handling the complexities of autonomous driving environments.
-For quantitative evaluation of our method’s semantic understanding capabilities, we generated segmentation maps for the images in the evaluation set using a pre-trained Segformer model [29] on the Cityscapes dataset [30].
-This semantic information allows us to assess semantic accuracy.
+For quantitative evaluation of our method’s semantic understanding capabilities, we generate segmentation maps for the images in the evaluation set using a pre-trained Segformer model [35] on the Cityscapes dataset [6].
+This semantic information allows us to assess the semantic accuracy of our proposed method.
+4.1.2 Text Dataset
+We generate a set of helping positives and canonical phrases for each query object using GPT-3.5 Turbo [25] for fine-tuning.
+For each primary query, the model provided 1 to 3 helping positive words and up to 6 canonical phrases.
+Figure 3 denotes the structure of one example for scene 12 from WayveScene101 in the generated dataset.
+The “Scene” description could be provided by the user or using an image captioning model.
+4.2 LLMs
+In addition to GPT-3.5 Turbo, we employ a variety of smaller, open-source models to improve inference efficiency and support on-device computation for applications such as autonomous vehicles and robotics.
+Specifically, we utilize instruction-tuned Qwen2.5 [37, 29], available in parameter sizes of 0.5, 1.5, 3, and 7 billion, along with instruction-tuned Llama 3.2 [10], offered in 1 and 3 billion parameter configurations.
+To optimize performance further, we fine-tune these models on a custom text dataset that we generated from GPT-3.5 Turbo.
+4.2.1 Fine-tuning Process
+We fine-tune the instruction-tuned Qwen and Llama models using LoRA [13] (Low-Rank Adaptation) for efficient parameter updates.
+We leverage Unsloth [7], which optimizes LoRA’s backward propagation, reducing FLOPs during gradient descent.
+The key projection layers (q_proj, k_proj, v_proj, o_proj) were targeted with a rank of 16 and a scaling factor of 16.
+The training is performed with a batch size of 2, gradient accumulation of 4 steps, and a learning rate of $1\times 10^{-4}$ using the AdamW optimizer and capped at 25 epochs.
+Mixed precision (FP16/BF16) accelerates training, while regular evaluations and early stopping ensure convergence.
+This optimized approach efficiently trains the models for generating canonical phrases and helping positives.
+(a) Qwen 0.5B Instruct (top) and fine-tuned (bottom)
+(b) Qwen 1.5B Instruct (top) and fine-tuned (bottom)
+(c) Qwen 3B Instruct (top) and fine-tuned (bottom)
+(d) Qwen 7B Instruct (top), fine-tuned (bottom)
+(e) Llama 1B (top) and fine-tuned (bottom)
+(f) Llama 3B (top) and fine-tuned (bottom)
+Figure 4: Qualitative comparison of semantic segmentation results before and after fine-tuning for different model variants.
+Each subfigure shows three views of the same scene, with the top row representing results from the instruction-tuned model (before fine-tuning) and the bottom row showing results from the fine-tuned model.
 4.3 Implementation
-We trained the LE3DGS using the same loss function as described in [21] on the Wayvesene101 dataset.
-We set $\lambda_{s}=0.5$, $\lambda_{CE}=\lambda_{smo}=\lambda_{u}=0.1$ and other weights to the deafault.
+We train the LE3DGS using the same loss function as described in [28] on the Wayvesene101 dataset.
+We set $\lambda_{s}=0.5$, $\lambda_{CE}=\lambda_{smo}=\lambda_{u}=0.1$ and other weights to the default.
 The training process involved two main phases: dense semantic feature extraction and 3D Gaussian point optimization.
@@
 The mean peak signal-to-noise ratio (PSNR) achieved is around 25.14.
-Table 1: Input structure for GPT-3.5 Turbo query generation in our implementation. This setup guides the language model to provide structured information about objects relevant to driving scenarios.
-Role Content
-System A helpful assistant for a computer vision task, providing structured information about objects to pay attention to while driving.
-Include both important objects and nearby objects that might be at the borders.
-Instructions for formatting the response, including: Main Positive (object to pay attention to) 1-4 Helping Positives (related terms or attributes) 4-6 Negatives (objects to differentiate from, including similar objects, nearby objects, and background elements)
-User "Driving through an intersection in an {Road Type} on a {Weather} {Time of Day}, what objects should the driver pay attention to? " or show the {Object}. "
-For the inference phase, we leveraged the API of GPT-3.5 Turbo [31] to dynamically generate queries.
-For each primary query, the model also provided 1 to 4 helping positive words and 4 to 6 canonical phrases.
-Table 1 denotes the input structure to the LLM model.
-The {Road Type}, {Weather}, and {Time of Day} in the user prompts could be provided from the dataset scene metadata [28].
+For the inference phase, we leveraged the API of GPT-3.5 Turbo [25] and the small Instruct-tuned and fine-tuned Qwen and Llama models running on the device to generate queries, helping positives and canonical words dynamically.
 We then calculated the relevancy score using Algorithm 1.
-To obtain object segmentations, we applied a threshold of 0.5 to the computed relevancy scores.
-Figure 1 illustrates our inference approach pipeline for scene 88 in Wayvescene101 dataset.
+We applied a threshold of 0.5 to the computed relevancy scores to obtain object segmentations.
 The canonical phrases used in the baseline approach are “object”, “things”, “stuff”, and “texture”, consistent with the approach described in [28].
+4.4 Results
+We evaluated the performance of Qwen and Llama model variants and GPT-3.5 Turbo against the baseline approach for open vocabulary segmentation through Intersection over Union (IoU), Accuracy, Precision, and mean Average Precision (mAP) metrics.
+Our experimental results, presented in Table 1, demonstrate that GPT-3.5 Turbo significantly outperforms the baseline approach across all metrics, achieving an IoU of $0.27$, accuracy of $0.90$, and precision of $0.37$ without helping positives, compared to the baseline’s IoU of $0.18$, accuracy of $0.64$, and precision of $0.19$.
+This substantial improvement validates the effectiveness of leveraging large language models for open vocabulary segmentation tasks.
+The fine-tuned variants of both Qwen and Llama models show marked improvements over their instruction-tuned counterparts.
+Notably, the Qwen2.5-1.5B-Finetuned model achieves the highest IoU ($0.28$) and precision ($0.37$) among all models in the configuration without helping positives, matching GPT-3.5 Turbo’s performance.
+Similarly, the Llama3.2-3B-Finetuned model demonstrates strong performance with an IoU of $0.26$ and accuracy of $0.84$, suggesting that effective fine-tuning can elevate the performance of smaller models to the levels approaching that of much larger models.
+Figure 4 shows the qualitative comparison of fine-tuned and instruct-tuned models.
+In our ablation study examining the quality of generated helping positives ($P_{help-pos}$), we observe that a model’s capacity significantly affects its ability to generate effective helping words.
+Larger models generate more semantically meaningful helping positives, leading to stable or improved segmentation performance, while smaller variants generate less effective helping words that can degrade segmentation quality.
+For instance, as shown in Table 1, the helping positives generated by Qwen2.5-7B-Finetuned enable robust segmentation performance (IoU: $0.23$, Accuracy: $0.89$, Precision: $0.34$), while helping positives from smaller models lead to more significant performance drops in the segmentation pipeline.
 Reconstructed
 Predefined $p_{\texttt{canon}}$
-Ours w/o $p_{\texttt{help}}$
-Ours
+GPT without $p_{\texttt{help-pos}}$
+GPT with $p_{\texttt{help-pos}}$
 Figure 5: Visual comparison of query relevance across different scenes with different approaches. From top to bottom query words are: “trees”, “traffic signs”, “pedestrian”, “sidewalk”, and “cars”.
-4.3 Results
-Our approach demonstrates significant improvements in open-vocabulary segmentation across various urban driving scenarios.
-Figure 2 provides a visual comparison, showcasing the progression from overly broad segmentation using predefined canonical phrases to refined results with our full LLM-enhanced method.
-Our approach accurately identifies and segments various elements such as trees, traffic lights, and sidewalks with high precision, illustrating the benefits of LLM-generated context-aware queries and helping positives.
-Inference Method $\mathrm{Accuracy}\uparrow$ $\mathrm{Presicion}\uparrow$ $\mathrm{mIoU}\uparrow$ $\mathrm{mAP}\uparrow$
-Predefined Canonical Phrase 0.65 0.33 0.23 0.49
-Ours w/o $p_{\texttt{help}}$ 0.84 0.42 0.33 0.53
-Ours 0.92 0.58 0.37 0.54
-Table 2: Comparison of inference methods: our approach (with and without helping positive words $p_{\texttt{help}}$) versus predefined canonical phrases.
-Quantitative results in Table 2 support these visual observations.
-Our full method achieves the highest performance across all metrics, with notable improvements in Accuracy (0.92), Precision (0.58), and mean Intersection over Union (mIoU) (0.37).
-These represent significant gains over the predefined canonical phrase method (0.65, 0.33, and 0.23 respectively).
-The ablation study (Ours w/o $p_{\texttt{help}}$) indicates the robustness of our LLM-driven query generation, outperforming the baseline even without helping positive words.
-(a) Mean Accuracy
-(b) Mean IoU
-(c) Mean Precision
-Figure 3: Performance comparison of our methods against baseline across different object classes for the scenes in dataset.
-Figure 3 further breaks down performance across various urban scene elements.
-Our method shows significant improvements in Mean IoU and Mean Precision for challenging classes like "sidewalk", "traffic light", and "traffic signs", while maintaining an edge in common classes such as "building" and "car".
-These results collectively demonstrate the effectiveness of our LLM-enhanced approach in improving scene understanding and object segmentation across diverse urban environments.
-For more detailed qualitative results, please refer to section A.1 in the appendix.
-5 Discussion
-Our approach demonstrates significant improvements in open-vocabulary 3D scene understanding, but there are several important aspects to consider:
-Limitations of 3D Gaussian Splatting
-While 3D Gaussian Splatting is effective for certain scenes, it struggles with generating large and dynamic scenes [32].
-This limitation can affect the overall performance and applicability of our approach in more dynamic environments, such as busy urban intersections or highways with fast-moving vehicles.
-System Integration and Real-time Performance
-The integration of our approach with existing autonomous driving systems presents both opportunities and challenges.
-While our method offers enhanced semantic understanding, its real-time performance in the context of rapidly changing driving scenarios requires further optimization.
-Current autonomous systems often rely on faster, albeit less semantically rich, perception methods.
-Bridging this gap involves not only improving our method’s computational efficiency but also developing strategies for seamless integration with existing sensor fusion and decision-making pipelines.
+The quality of helping positives reaches its peak with GPT-3.5 Turbo, whose generated helping words enable substantial improvements in segmentation performance, achieving IoU ($0.32$, $+18.5\%$), accuracy ($0.94$, $+4.4\%$), and precision ($0.50$, $+35.1\%$).
+This marked enhancement suggests that larger-scale models can generate more contextually appropriate and semantically rich helping positives, leading to more effective scene segmentation.
+The quantitative results are further supported by qualitative comparisons shown in Figure 5, which demonstrate improved segmentation quality across various urban scene elements.
 5 Conclusion
-In this study, we proposed a novel method that integrates LLMs with LE3DGS to improve autonomous vehicle scene understanding.
-By leveraging the robust reasoning capabilities of LLMs and the detailed spatial representation of 3D Gaussians, our approach significantly improves the accuracy and precision of open vocabulary object segmentation and scene interpretation tasks.
-Our experimental results indicate that our method outperforms traditional predefined canonical phrase-based approaches on all aspects of accuracy, precision, and mean Intersection over Union.
-Future work will focus on enhancing our system’s capability to handle dynamic scenes by integrating LLMs with advanced 3D Gaussian Splatting techniques.
-We aim to fuse our approach with multi-modal sensor data, such as LiDAR, to create a more robust and comprehensive perception system.
-Another crucial next step is to seamlessly integrate our advanced semantic understanding with autonomous driving planners.
-This integration will involve developing efficient interfaces that translate our rich semantic insights into actionable information, directly informing path planning and decision-making algorithms for safer and more intelligent autonomous navigation.
-Appendix A Appendix
-A.1 Qualitative Results
-Figures 4, 5, and 6 showcase our approach’s effectiveness in open vocabulary segmentation across diverse driving scenes.
-Each figure presents, from top to bottom: ground truth images, Segformer-generated segmentation maps (used for quantitative evaluation), our high-quality scene reconstructions using LE3DGS, and our open vocabulary segmentation results.
-Our method accurately highlights key elements like cars, pedestrians, trees, buildings, and traffic signs/lights.
-For quantitative evaluation, we align with Segformer classes [33], but our visualizations include additional classes to demonstrate the flexibility of our approach.
-These figures illustrate our method’s capability to accurately reconstruct and semantically segment complex urban scenes, even for classes beyond traditional segmentation models.
-Figure 4: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scene 88.
-(a) Scene 12
-(b) Scene 21
-Figure 5: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scenes 12 and 21.
-(a) Scene 36
-(b) Scene 78
-Figure 6: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scenes 36 and 78.
+In this study, we proposed a novel method that integrates LLMs with LE3DGS to enhance scene understanding for autonomous vehicles.
+Our approach significantly improves open vocabulary object segmentation in complex driving environments, substantially outperforming traditional predefined canonical phrase-based approaches.
+Our results demonstrate that carefully fine-tuned smaller language models can achieve performance comparable to larger expert models while enabling on-device applications.
+Furthermore, our ablation studies revealed that the effectiveness of helping positives correlates with model scale, with larger models better equipped to leverage additional semantic information.
+Future work will focus on integrating this semantic understanding with autonomous driving planners, developing efficient interfaces that translate semantic insights into actionable information for path planning and decision-making algorithms.